OpenStudy (anonymous):

what's the sqrt of 605x^3y^4

OpenStudy (anonymous):

\[\sqrt{605x^3y^4} = (605x^3y^4)^\frac{ 1 }{ 2 } = \sqrt{605}x^\frac{ 3 }{ 2 }y^2\]

OpenStudy (anonymous):

thank-you

OpenStudy (mathstudent55):

You need to find the largest perfect squares inside the square root, and take them out: \( \sqrt{605x^3y^4} \) \( = \sqrt{5(11^2)x^2x(y^2)^2} \) \( = 11xy^2\sqrt{5x} \)